### v3

[…]
Jiang, Ye
[…]
Sample Complexity for Constrained Markov Decision Process
[…]
Achieving Instance-dependent Sample Complexity for Constrained Markov Decision Process
[…]
$\text{Jiashuo Jiang}^{{\dagger}}$, $\leavevmode\nobreak\ \text{Yinyu Ye}^{{\ddagger}\S}$
[…]
We consider the reinforcement learning problem for the constrained Markov decision process (CMDP), which plays a central role in satisfying safety or resource constraints in sequential learning and decision-making. In this problem, we are given finite resources and a MDP with unknown transition probabilities. At each stage, we take an action, collecting a reward and consuming some resources, all assumed to be unknown and need to be learned over time. In this work, we take the first step towards deriving optimal problem-dependent guarantees for the CMDP problems. We derive a logarithmic regret bound, which translates into a $O(\frac{1}{\Delta\cdot\varepsilon}\cdot\log^{2}(1/\varepsilon))$ sample complexity bound, with $\Delta$ being a problem-dependent parameter, yet independent of $\varepsilon$. Our sample complexity bound is the first problem-dependent guarantee for CMDP problems. To achieve this advance, we develop a new framework for analyzing CMDP problems. To be specific, our algorithm operates in the primal space and we resolve the primal LP for the CMDP problem at each period in an online manner, with adaptive remaining resource capacities. The key elements of our algorithm are: i) a characterization of the instance hardness via LP basis, ii) an eliminating procedure that identifies one optimal basis of the primal LP, and; iii) a resolving procedure that is adaptive to the remaining resources and sticks to the characterized optimal basis.
[…]
Constrained MDP, Reinforcement Learning, Online Packing/Linear Programming, Problem-dependent Bound
1 Introduction
Reinforcement learning (RL) is pivotal in the realm of dynamic decision-making under uncertainty, where the objective is to maximize total reward through ongoing interaction with and learning from an enigmatic environment. Markov Decision Processes (MDPs) are a prevalent framework for encapsulating environmental dynamics. MDPs have been instrumental in various domains, such as inventory management (Alvo et al. 2023), video gaming (Mnih et al. 2013), robotics (Kober et al. 2013), recommender systems (Shani et al. 2005), and beyond. Yet, they fall short in accommodating additional constraints that may influence the formulation of the optimal policy and the decision-maker’s engagement with the uncertain environment. Often, in MDP applications, there are stringent constraints on utilities or costs, emanating from areas like safe autonomous driving (Fisac et al. 2018), robotics (Ono et al. 2015), and financial management (Tamar et al. 2012). These constraints might also symbolize limitations on resources within resource allocation contexts with applications to revenue management (Jia et al. 2023, Jiang 2023, Li et al. 2023), inventory management and queueing systems (Chen et al. 2021). Constrained MDPs (CMDPs), as introduced in Altman (1999), enhance MDPs to factor in constraints affecting long-term policy results. In CMDPs, the decision-maker aims to optimize cumulative rewards while adhering to these constraints. Our paper focuses on CMDPs, and we aim to develop efficient algorithmic solutions.
The significance of RL in CMDP contexts has garnered substantial attention in recent years. A variety of methods for tackling CMDPs have been developed, including the primal-dual technique (e.g. Efroni et al. (2020)) which leverages the Lagrangian dual of CMDPs and implements an online learning strategy for the iterative update of dual variables. Other methods encompass constrained optimization (Achiam et al. 2017, Sun et al. 2024), the Lyapunov technique (Chow et al. 2018), among others. Previous research has established minimax bounds for CMDPs, representing the optimal regret that can be achieved for the most challenging problem within a specific class of problems. Nonetheless, these minimax regret bounds and worst-case scenarios can be overly conservative, leading to a disconnect between theoretical guarantees and practical performance for specific problem instances. A more tailored approach is warranted—one that ensures great performance on every single problem instance and furnishes problem-dependent guarantees. Our research takes the first step towards deriving optimal problem-dependent guarantees for CMDP problems.
1.1 Main Results and Contributions
In the exploration of the $\gamma$-discounted infinite horizon CMDP problem within a tabular context, we leverage a generative model that (as studied in Kearns et al. (2002), Kakade (2003), Li et al. (2024)), for each state-action pair, yields a sample from the unknown stochastic reward, costs, and state transitions. The main result of our research is the introduction of a novel algorithm that promises a $O(\kappa\cdot\log N)$ regret upper bound for both the cumulative reward and constraint violations, where $N$ is the number of queries to the generative model and $\kappa$ is a problem-specific parameter, yet independent of $N$. This translates into a $\tilde{O}\left(1/\varepsilon\right)$ sample complexity bound, where $\tilde{O}(\cdot)$ hides a term polynomial in $\log(1/\varepsilon)$. Note that the state-of-the-art sample complexity bound under the worst-case scenarios is $O(1/\varepsilon^{2})$, which has been established in a series of work (Wang 2020, Jin and Sidford 2020, Efroni et al. 2020). Though the $O(1/\varepsilon)$ or better iteration complexity has been achieved in Li et al. (2021b), Liu et al. (2021b), Zhou et al. (2022), Gladin et al. (2023), it comes with a sample complexity bound no better than $O(1/\varepsilon^{2})$. Our algorithm enjoys a sample complexity bound that has a better dependency in terms of $\varepsilon$. To achieve these problem-dependent guarantees, we develop several elements.
First, we develop new characterizations of the problem instance hardness for CMDP problems. Note that a key component for achieving instance-dependent bounds is to characterize the “hardness” of the underlying problem instance. That is, we need to identify a positive gap to separate the optimal policy from the sub-optimal policies for a particular problem instance. The importance of identifying such a gap has been demonstrated in instance-optimal learning for multi-arm-bandit problems (e.g. Lai and Robbins (1985)) and reinforcement learning problems (e.g. Auer et al. (2008)), among others. The gap characterizes the hardness of the underlying problem instance (e.g. Simchowitz and Jamieson (2019)). For CMDP, identifying such a gap is non-trivial because the optimal policies for CMDP are randomized policies (Altman 1999). Then, the policies can be represented by distributions over the action set and the sub-optimal policies can be arbitrarily close to the optimal policy as long as the corresponding sub-optimal distributions converge to the optimal one. To tackle this problem, we show that the feasible region for the policies can be represented as a polytope and we only need to focus on the corner points of this polytope to find an optimal policy. Therefore, the hardness can simply be characterized as the distance between the optimal corner point and the sub-optimal corner point, as illustrated in detail in Section 3.1. This is the first characterization of problem instance hardness for CMDP problems.
Then, we devise a new algorithmic framework to analyze CMDP problems, inspired by the online packing/linear programming (LP) literature (Agrawal et al. 2014, Kesselheim et al. 2014, Li and Ye 2022, Li et al. 2021c). Specifically, we utilize a linear programming reformulation of the CMDP problem, where policies are delineated via occupancy measures (Altman 1999). The optimal policy emerges from the LP’s solution; however, the indeterminate model parameters mean the LP cannot be solved directly but must be approached online as we obtain more samples from the generative model. Each generative model query leads to solving an empirical LP with accrued samples, and our final policy is derived from averaging these solutions—an approach akin to the methodology in online LP. A critical feature of our algorithm is the adaptiveness of the LP constraints’ right-hand side to the input samples, a technique demonstrated to achieve logarithmic regret in online LP literature, which we now apply to CMDP problems.
Furthermore, we extend our contributions to the online LP literature. The prevalent assumption of non-degeneracy in online LP, which postulates a unique LP solution, is not a requirement in our approach. We introduce an innovative method to identify and consistently apply one optimal LP basis, even when multiple bases exist. Aligning with a single optimal basis—as opposed to requiring uniqueness—affords the same benefits and leads to logarithmic regret. To delve into details, our algorithm enhances performance through two new components. The first is a systematic procedure to pinpoint a particular optimal LP basis from potentially several (outlined in Algorithm 1), diverging from the traditional literature that presupposes a sole optimal solution and basis. The second component is a resolving procedure that capitalizes on the structure of the identified optimal basis. Once an optimal basis is selected, we can determine the corresponding optimal solution by solving a set of linear equations, which involves only the non-zero basic variables and the active constraints. This resolving process transforms the algorithmic idea of resolving an LP, as seen in previous works (Agrawal et al. 2014, Bumpensanti and Wang 2020, Li and Ye 2022, Vera and Banerjee 2021, Li et al. 2021c), into solving specific linear equations based on the chosen optimal basis (detailed in Algorithm 2). Previous research relies on the non-degeneracy assumption to stabilize the optimal basis throughout the resolving process. However, our two innovative elements enable our algorithm to consistently adhere to one optimal basis, thus removing the non-degeneracy assumption. This not only simplifies the algorithm’s execution (from resolving LP to resolving linear equations) but also extends its applicability, freeing it from assumptions that have traditionally limited the scope of online LP techniques.
We finally discuss the dependency of our sample complexity bound on problem parameters other than $\varepsilon$. We restrict to the MDP context without resource constraints. Denote by $\mathcal{S}$ the state set and $\mathcal{A}$ the action set. We show a sample complexity bound of $O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right)$, where $\Delta$ is the constant that represents the hardness of the underlying problem instance. Compared to the optimal worst-case sample complexity $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$ that is achieved in a series of work (Sidford et al. 2018, Wainwright 2019, Wang 2020, Agarwal et al. 2020, He et al. 2021), our bound has a worse dependency over $|\mathcal{S}|$ and $1-\gamma$. This is due to our algorithm being LP-based and the dimension of the LP ($|\mathcal{S}|$ and $1-\gamma$) will influence our final bounds. However, our bound enjoys a better dependency in terms of $\varepsilon$. For the general CMDP problem, our bound will depend additionally on the conditional number of the constraint matrix in the LP formulation, which is a byproduct of the resolving LP heuristics (Vera and Banerjee 2021, Li et al. 2021c). However, our $\tilde{O}(1/\varepsilon)$ sample complexity bound depends polynomially on other parameters including $|\mathcal{S}|$, $|\mathcal{A}|$, $1-\gamma$, and the number of constraints. Partial results of this paper have been accepted at NeurIPS 2024 conference. Compared with the conference version, the current paper not only enriches the interpretations significantly and includes complete detailed proofs, but also includes the following key extensions.
Extension to the finite horizon episodic setting. We extend our approaches and results to the finite horizon setting, where there are multiple periods in a horizon and the horizon can repeated by multiple times to learn the optimal policy. Note that the optimal policy under the finite horizon setting can still be written as a LP. We directly extend the online LP approach to handle the finite horizon setting and obtain similar results. To be specific, we show that in order to obtain an $\epsilon$-approximate policy where we approximate the optimal reward by a $\varepsilon$ gap and allows at most a $\varepsilon$ constraint violation, we need at most $\tilde{O}(1/\varepsilon)$ number of samples. The details are given in Section 7.1.
Extension to the online and offline learning settings. We consider more involved sampling procedure beyond assuming generative model. Note that in practice, the agent usually interacts with the true environment (online learning setting), where the agent can only obtain the information regarding the state it is currently visiting and the action it is currently taking, or interacts with an offline simulator (offline learning setting), where there is a fixed historical data set which may be collected using some other algorithms. Different from the generative model where we can obtain information for arbitrary states and actions, the online and offline learning settings put further restrictions on the freedom of obtaining samples and require additional considerations when designing algorithms. However, we show that our approach can still be directly extended to handle the online and offline learning settings, and we are also able to derive the similar $\tilde{O}(1/\varepsilon)$ sample complexity. The details are presented in Section 7.2.
In summary, our work presents a new algorithm for CMDP problems to achieve the problem-dependent logarithmic regret or $\tilde{O}(1/\varepsilon)$ sample complexity bound, for the first time in the literature. Firstly, we have pioneered the application of the online LP framework to analyze CMDP problems. Our algorithm operates in the primal space, in contrast to the primal-dual framework associated with worst-case guarantees (e.g. Ding et al. (2021)). Secondly, our novel approach to online LP results in logarithmic regret without the non-degeneracy assumption, expanding its applicability and providing new insights into both CMDP and online LP problems. We finally explore extensions to the finite horizon setting, as well as the online and offline learning settings, and obtain similar $\tilde{O}(1/\varepsilon)$ sample complexity bounds. These extensions demonstrate the generality of our approach.
1.2 Further Related Work
Deriving the (logarithmic) problem-dependent guarantees is a long-standing topic in online decision-making (Lai and Robbins 1985, Graves and Lai 1997). Notably, there has been a recent trend of deriving (logarithmic) problem-dependent guarantees for reinforcement learning problems, where a series of work (Zanette and Brunskill 2019, Simchowitz and Jamieson 2019, Mou et al. 2020, Foster et al. 2020, Dann et al. 2021, Bhandari and Russo 2021, Wagenmaker et al. 2022, Wagenmaker and Foster 2023, Xu and Zeevi 2024, Duan and Wainwright 2024, Chen et al. 2024) has proposed various ways to characterize the hardness of the problem instance. A prevalent way to characterize the hardness is to define a constant that represents the policy sub-optimality gap between the optimal policy and a policy that takes a sub-optimal action given a state. Then, the logarithmic guarantees depend on this constant. Compared to the previous approaches and guarantees, our guarantees adopt this idea of policy sub-optimal gap. Moreover, our approach can handle resource or safety constraints that need to be satisfied by the policy. In this way, our approach generalizes to multi-objective (safe) RL problems. Safe reinforcement learning problems have been investigated extensively under various settings in a series of work (Liu et al. 2021a, Wei et al. 2021, Ding et al. 2021, Li et al. 2021b, Ghosh et al. 2022, Chen et al. 2022, Ding and Lavaei 2023, Germano et al. 2023, Gladin et al. 2023), where the $O(\sqrt{T})$ regret bounds or $O(1/\varepsilon^{2})$ sample complexity bounds are proved.
Our method is motivated by the online LP literature. The online LP problem (Molinaro and Ravi 2014, Agrawal et al. 2014, Gupta and Molinaro 2014) covers a wide range of applications through different ways of specifying the underlying LP, including secretary problem (Ferguson 1989), online knapsack problem (Arlotto and Xie 2020), quantity-based network revenue management problem (Jasin and Kumar 2012), network routing problem (Buchbinder and Naor 2009), matching problem (Mehta et al. 2007), etc. Notably, the problem has been studied under either (i) the stochastic input model where the coefficient in the objective function, together with the corresponding column in the constraint matrix is drawn from an unknown distribution $\mathcal{P}$, or (ii) the random permutation model where they arrive in a random order (Molinaro and Ravi 2014, Agrawal et al. 2014, Kesselheim et al. 2014, Gupta and Molinaro 2014). With an additional non-degeneracy assumption, the logarithmic regret has been derived for the quantity-based network revenue management problem (Jasin and Kumar 2012, Jasin 2014), the online LP problem (Li and Ye 2022), and the more general convex allocation problem (Ma et al. 2022). The non-degeneracy assumption has been relaxed in a series of work (Bumpensanti and Wang 2020, Vera and Banerjee 2021, Jiang et al. 2022, Wei et al. 2023, Ao et al. 2024, Banerjee and Freund 2024) under various settings.
1.3 Roadmap
In Section 2, we describe the formal problem formulation and introduce the notations that we use in the paper. In Section 3, we present the linear programming reformulation of our problem and we give a general overview of our approach to illustrate the main idea. Then, in Section 4, we describe how to construct the estimates of the LP reformulation of the CMDP problems, which will be used in our algorithm. The formal algorithm is presented in Section 5 and the corresponding theoretical analysis is conducted in Section 6.
2 Preliminaries
We consider a CMDP problem with a finite set of states $\mathcal{S}=\{1,2,\dots,|\mathcal{S}|\}$ and a finite set of actions $\mathcal{A}=\{1,2,\dots,|\mathcal{A}|\}$. We denote by $\gamma\in(0,1)$ a discount factor. We also denote by $P:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}(\mathcal{S})$ the probability transition kernel of the CMDP, where $\mathcal{D}(\mathcal{S})$ denotes a probability measure over the state space $\mathcal{S}$. Then, $P(s^{\prime}|s,a)$ denotes the probability of transiting from state $s$ to state $s^{\prime}$ when the action $a$ is executed. The initial distribution over the states of the CMDP is denoted by $\mu_{1}$.
There is a stochaastic reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}[0,1]$ and $K$ stochastic cost functions $c_{k}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}[0,1]$ for each $k\in[K]$. We also denote by $\hat{r}(s,a)=\mathbb{E}[r(s,a)]$ for each $(s,a)$ and $\hat{c}_{k}(s,a)=\mathbb{E}[c_{k}(s,a)]$ for each $(s,a)$, for each $k$. For any Markovian policy $\pi$, where the action of $\pi$ depends only on the current state and the action of $\pi$ is allowed to be randomized, we denote by $V_{r}(\pi,\mu_{1})$ the infinite horizon discounted reward of the policy $\pi$, with the formulation of $V_{r}(\pi,\mu_{1})$ given below:
$V_{r}(\pi,\mu_{0})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot r(s_{t},a_{t})\mid\mu_{1}\right],$ (1)
where $(s_{t},a_{t})$ is generated according to the policy $\pi$ and the transition kernel $P$ with the initial state distribution $\mu_{1}$. For each $k\in[K]$, the infinite horizon discounted cost of the policy $\pi$ is denoted by $V_{k}(\pi,\mu_{1})$, and the following constraint needs to be satisfied by the policy $\pi$,
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\leq\alpha_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\in[K].$ (2)
To solve the CMDP problem, we aim to find an optimal Markovian policy, denoted by $\pi^{*}$, that maximizes the reward in (1) while satisfying the cost constraint (2) for each $k\in[K]$, with $\alpha_{k}\in\left[0,\frac{1}{1-\gamma}\right]$ being a pre-specified value for each $k\in[K]$. Importantly, we assume that the reward function $r$, the cost functions $\{c_{k}\}_{k=1}^{K}$, and the transition kernel $P$, are all unknown to the decision maker. Our goal is to obtain a policy $\pi$ that approximates the optimal policy $\pi^{*}$ with as few samples as possible. We now describe the sampling procedure and present the performance measure of our policy.
We assume the existence of a stylized generative model $\mathcal{M}$. The model $\mathcal{M}$ satisfies the following condition. {assumption} For each state and action pair $(s,a)$, we can query the model $\mathcal{M}$ to obtain an observation of $r(s,a)$, $c_{k}(s,a)$ for each $k\in[K]$, and the new state $s^{\prime}\in\mathcal{S}$, where the transition from $s$ to $s^{\prime}$ follows the probability kernel $P(s^{\prime}|s,a)$ independently.
Note that in reinforcement learning for CMDP problems, querying the generative model $\mathcal{M}$ can be costly. Therefore, it is desirable to query the model $\mathcal{M}$ as less as possible, while guaranteeing the near optimality of the approximate policy. To this end, we study two measures of the performance of our policy. One is regret, i.e., conditioning on we can query the model $\mathcal{M}$ at most $N$ times, we aim to derive a policy $\pi$ and bound
$\text{Regret}_{r}(\pi,N):=V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\text{Regret}_{k}(\pi,N):=V_{k}(\pi,\mu_{1})-\alpha_{k},\leavevmode\nobreak\ \forall k\in[K].$ (3)
The other one is sample complexity, i.e., for any $\varepsilon$, we aim to find an $\varepsilon$-accurate policy $\pi$ such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\pi,\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K],$ (4)
with as few samples as possible. Indeed, these two measures can be transferred to each other, and in our analysis, we derive the sample complexity bound from the regret bound.
Remark. Note that there are other formulations of the constraints in CMDP, for example,
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\geq\lambda_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\in[K].$ (5)
in a series of work that studies safe reinforcement learning. However, the formulation (5) can be transferred from the formulation (2). One can set $\alpha_{k}=\frac{1}{1-\gamma}-\lambda_{k}$ for each $k\in[K]$, and it is easy to see that the two inequalities are equivalent to each other,
$\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\geq\lambda_{k}\Leftrightarrow\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot(1-c_{k}(s_{t},a_{t}))\mid\mu_{1}\right]\leq\alpha_{k}.$
Therefore, we can equivalently use formulation (2) with the cost function defined as $1-c_{k}$ for each $k\in[K]$.
2.1 Notations
Throughout the paper, we use $n\in[N]$ to denote the time step. For the infinite horizon discounted problem in Section 2, we can query the generative model $\mathcal{M}$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ for one time, for a time step $n\in[N]$.
For a matrix $A$, we denote by $\mathcal{J}$ a row index set and $\mathcal{I}$ a column index set. Then, $A(:,\mathcal{I})$ denotes a sub-matrix of $A$ with columns drawn from the index set $\mathcal{I}$, and $A(\mathcal{J},:)$ denotes a sub-matrix of $A$ with rows drawn from the index set $\mathcal{J}$. Similarly, $A(\mathcal{J},\mathcal{I})$ denotes a sub-matrix of $A$ with rows drawn from $\mathcal{J}$ and columns drawn from $\mathcal{I}$. For any vector $\bm{\beta}$, we use $\bm{\beta}_{\mathcal{J}}$ to denote a sub-vector of $\bm{\beta}$ with element drawn from the index set $\mathcal{J}$.
3 Problem Reformulation and Overview of Our Approach
The infinite horizon discounted setting described in Section 2 admits a linear programming reformulation. To be specific, due to the existence of the constraints, the optimal policy of a CMDP can be randomized policies (e.g. Altman (1999)), where it is optimal to take a stochastic action given the current state. Therefore, it is convenient to represent a policy through the occupation measure, which gives us the desired linear programming reformulations of the CMDP problems.
For the infinite horizon discounted problem, the occupancy measure is defined as $q_{\pi}(s,a)$ for any state $s$, action $a$, and policy $\pi$. Note that $q_{\pi}(s,a)$ represents the total expected discounted time spent on the state-action pair $(s,a)$, under policy $\pi$, multiplied by $1-\gamma$. Then, following Altman (1999), the optimal policy (and the optimal occupancy measure) can be obtained from the following linear programming.
$V^{\mathrm{Infi}}=\leavevmode\nobreak\ \max$ $\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{r}(s,a)\cdot q(s,a)$ (6a)
$\mathrm{s.t.\ }$ $\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{c}_{k}(s,a)\cdot q(s,a)\leq\alpha_{k}$ $\forall k\in[K]$ (6b)
$\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}q(s^{\prime},a)\cdot(\delta_{s,s^{\prime}}-\gamma\cdot P(s|s^{\prime},a))=(1-\gamma)\cdot\mu_{1}(s)$ $\forall s\in\mathcal{S}$ (6c)
$q(s,a)\geq 0$ $\forall s\in\mathcal{S},a\in\mathcal{A},$ (6d)
where $\delta_{s,s^{\prime}}=\mathbbm{1}_{s=s^{\prime}}$, and $\mu_{1}(s)$ denotes the probability for the first state to be realized as $s\in\mathcal{S}$ following the initial distribution $\mu_{1}$. Note that for an optimal solution $\{q^{*}(s,a)\}_{\forall s\in\mathcal{S},\forall a\in\mathcal{A}}$ to (6), the corresponding optimal policy $\pi^{*}$ will be
$P(\pi^{*}(s)=a)=\left\{\begin{aligned} &\frac{q^{*}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})},&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})=0,\end{aligned}\right.$ (7)
where $P(\pi^{*}(s)=a)$ denotes the probability for the policy $\pi^{*}$ to take the action $a\in\mathcal{A}$ given the state $s\in\mathcal{S}$. In fact, when $\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})=0$, we can take an arbitrary action. In what follows, we rely on the linear programming formulation (6) to derive our results.
3.1 Characterization of Instance Hardness
We illustrate our approach to solving the CMDP problems and present the formal algorithms. Our approach is based on the linear programming reformulations presented in (6). We now rewrite the LP (6) into the following standard formulation to proceed with our illustration.
$V=$ $\max$ $\hat{\bm{r}}^{\top}\bm{q}$ (8)
    s.t. $C\bm{q}\leq\bm{\alpha}$
$B\bm{q}=\bm{\mu}$
$\bm{q}\geq 0.$
For the infinite horizon discounted problem described in Section 2, the matrix $B\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$ and the element at $s$-th row and $(s^{\prime},a^{\prime})$th column of the matrix $B$ is $(\delta_{s,s^{\prime}}-\gamma\cdot P(s|s^{\prime},a^{\prime}))$. The matrix $C\in\mathbb{R}^{K\times|\mathcal{S}|}$, and the element at the $k$-th row and $(s,a)$-th column of the matrix $C$ is $\hat{c}_{k}(s,a)$. The vector $\bm{\mu}\in\mathbb{R}^{|\mathcal{S}|}$ and the $s$-th element of $\bm{\mu}$ is $(1-\gamma)\cdot\mu_{1}(s)$, for $s\in\mathcal{S}$.
Note that one crucial step for achieving problem-dependent bounds is to characterize the hardness of the underlying problem instance and define a gap that separates the optimal policies from the others. For multi-arm-bandit problem, the characterization of the hardness can be the gap between the optimal arm and the best sub-optimal arm (e.g. Lai and Robbins (1985)). For reinforcement learning problem, the characterization of the hardness can also be the gap between the optimal policy and the best sub-optimal policy (e.g. Auer et al. (2008)). As long as this separation gap is a positive constant, denoted by $\Delta$, separating the optimal policy from the others with a probability at least $1-\epsilon$ would require samples at most $\frac{1}{\Delta}\cdot\log(1/\epsilon)$, which finally implies the instance-optimal sample complexity bound.
For CMDP problem, characterizing the hardness of the problem instance can be hard. Based on LP (8), we know that a feasible policy corresponds to a feasible solution and the sub-optimal solution can be arbitrarily close to the optimal solution since the feasible set is “continuous”. Therefore, there is no direct way to identify a positive gap between the optimal policies and the sub-optimal ones. However, from standard LP theory, we know that one corner point of the feasible region must be one optimal solution. Therefore, we can simply focus on the corner points when solving LP (8) and we define the gap as the distance between the optimal corner point and the sub-optimal corner point, as illustrated in Figure 1. As we will show later, the LP reformulation (8) and such a characterization of hardness via corner points will inspire our entire approach.
Figure 1: A graph illustration of the hardness characterization via LP basis, where the shaded area denotes the feasible region for the policies.
Since the problem hardness is characterized via corner points, it is essential to provide further characterization of the corner points. Note that in LP theory, the corner point is called basic solution and can be represented by LP basis, which involves the set of basic variables that are allowed to be non-zero, and the set of active constraints that are binding under the corresponding basic solution. For our approach, we identify one optimal basic solution and then stick to this optimal basic solution. However, there can be multiple optimal bases of the LP (8). We characterize one optimal basis and we exploit its structure during the entire horizon. We first observe the following result regarding an optimal basis to the LP (8), which follows from standard LP theory. The proof of Lemma 3.1 is relegated to Section 9 for completeness.
Lemma 3.1
Denote by $b$ the number of rows in the matrix $B$. Then, there exists a subset $\mathcal{J}^{*}\subset[K]$ with $K^{\prime}=|\mathcal{J}^{*}|$ and an optimal solution $\bm{q}^{*}$ to LP (8) such that there are $b+K^{\prime}$ variables in $\bm{q}^{*}$ are non-zero. Moreover, we denote by $\mathcal{I}^{*}$ the index set of the non-zero element in $\bm{q}^{*}$. Then, the optimal solution $\bm{q}^{*}$ can be uniquely determined as the solution to the linear system
$C(\mathcal{J}^{*},\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\alpha}_{\mathcal{J}^{*}},$ (9a)
$B(:,\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\mu},$ (9b)
$\bm{q}_{\mathcal{I}^{*c}}=0.$ (9c)
with $\mathcal{I}^{*c}$ being the complementary set of the index set $\mathcal{I}^{*}$.
3.2 Overview of Our Approach
In the first part, we aim to identify the set $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, which characterizes one optimal basis of LP (6). By doing so, we also obtain the optimal actions to take given each state (and each period), by noting the relationship between the optimal actions and the optimal solution (occupancy measure) in (7). In the second part, we aim to obtain the optimal solution (occupancy measure) by solving the linear equations (9). However, since the parameters of the linear equations (9) are unknown, we cannot directly solve (9). Instead, we need to use historical samples to construct estimates of the parameters of (9) and gradually approximate the optimal solution. To this end, we seek tools from the online packing/linear programming literature (e.g. Agrawal et al. (2014), Li and Ye (2022), Vera and Banerjee (2021), Bumpensanti and Wang (2020)), where we interpret the right-hand side (RHS) of the linear equations (9) as resource capacities. Note that a crucial element in deriving the logarithmic or better problem-dependent regret in the previous literature is to let our algorithm be adaptive to the remaining resource capacities in real-time. We adopt this adaptive resolving idea in our algorithm.
In the remaining sections, we first describe how our LP-based approach implies a characterization of the “hardness” of the underlying problem instance. We then describe how to construct estimates of unknown parameters for the CMDP problem and we present our formal algorithms. We show how to identify one optimal basis, and how to approximate the optimal solution with an adaptive constraint design.
4 Construct Estimates in CMDP
In this section, we describe how to construct estimates for LP (8). To this end, for a round $N_{0}$, we denote by $\mathcal{F}_{N_{0}}$ the filtration of all the information collected up to round $N_{0}$. Then, we denote by $\bar{C}_{N_{0}}$ (resp. $\bar{B}_{N_{0}}$) an estimate of the matrix $C$ (resp. $B$), constructed using the information in the set $\mathcal{F}_{N_{0}}$. We also denote by $\bar{\bm{r}}_{N_{0}}$ and estimate of $\hat{\bm{r}}$ constructed from the information in the set $\mathcal{F}_{N_{0}}$. Then, similar to Efroni et al. (2020), we can use the following LP to obtain an estimate of $V$ (8).
$\bar{V}_{N_{0}}=$ $\max$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ (10)
    s.t. $\bar{C}_{N_{0}}\bm{q}\leq\bm{\alpha}+\lambda_{N_{0}}$
$\bar{B}_{N_{0}}\bm{q}\leq\bm{\mu}+\lambda_{N_{0}}$
$\bar{B}_{N_{0}}\bm{q}\geq\bm{\mu}-\lambda_{N_{0}}$
$\bm{q}\geq 0,$
with $\lambda_{N_{0}}$ being a parameter that we specify later. To bound the estimation gap, it is useful to bound the optimal dual solution to (8). To this end, we adopt the approach in Hiriart-Urruty and Lemaréchal (1996), Nedić and Ozdaglar (2009) that utilizes Slater’s condition, which is imposed as an assumption below. {assumption} There exists a policy $\bar{\pi}$ such that all the resource constraints are satisfied strictly. In other words, there exists an occupancy measure $\bar{\bm{q}}$ such that $B\bar{\bm{q}}=\bm{\mu}$ and $C\bar{\bm{q}}<\bm{\alpha}$. In fact, for each state $s\in\mathcal{S}$, there exists a null action that consumes no resource. The Slater point $\bar{\bm{q}}$ can be set as the policy that takes the null action given each state. The estimation error will be related to the gap between the Slater point $\bar{\bm{q}}$ and the optimal point $\bm{q}^{*}$. We then define the lower gap as
$\mathrm{Gap}_{1}(N_{0},\varepsilon)\geq V-\bar{V}_{N_{0}}$ (11)
and the upper gap as
$\mathrm{Gap}_{2}(N_{0},\varepsilon)\geq\bar{V}_{N_{0}}-V$ (12)
with both inequalities (11) and (12) hold with probability at least $1-\varepsilon$.
4.1 Estimates for the Infinite Horizon Discounted Problem
We define the following empirical estimates over the LP parameters.
$\bar{r}_{N_{0}}(s,a)=\frac{\sum_{n=1}^{N_{0}}r^{n}(s,a)}{N_{0}},\bar{c}_{k,N_{0}}(s,a)=\frac{\sum_{n=1}^{N_{0}}c^{n}_{k}(s,a)}{N_{0}},\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }\bar{P}_{N_{0}}(s^{\prime}|s,a)=\frac{\sum_{n=1}^{N_{0}}\mathbbm{1}_{s^{n}(s,a)=s^{\prime}}}{N_{0}},$ (13)
where $r^{n}(s,a)$ denotes the $n$-th observation of the reward, and $c^{n}_{k}(s,a)$ denotes the $n$-th observation of the $k$-th cost, and $s^{n}(s,a)$ denotes the $n$-th observation of the state transition for the state-action pair $(s,a)$, for $n\in[N_{0}]$. We denote by
$\mathrm{Rad}(N_{0},\varepsilon)=\sqrt{\frac{\log(2/\varepsilon)}{2N_{0}}}.$
Following the standard Hoeffding’s inequality, we know that $|\bar{r}_{N_{0}}(s,a)-\hat{r}(s,a)|$, $|\bar{c}_{k,N_{0}}(s,a)-\hat{c}_{k}(s,a)|$, and $|\bar{P}_{N_{0}}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|$ are all upper bounded by $\mathrm{Rad}(N_{0},\varepsilon)$ with probability at least $1-\varepsilon$. Then, we use the following LP with empirical estimates to approximate the value of $V^{\mathrm{Infi}}$, as given in (10). Here, $\bar{C}_{N_{0}}\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{A}|}$, and the element at the $k$-th row and $(s,a)$-th column of the matrix $\bar{C}_{N_{0}}$ is $\bar{c}_{k,N_{0}}(s,a)$. $\bar{B}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$, and the element at $s$-th row and $(s^{\prime},a^{\prime})$-th column of the matrix $\bar{B}_{N_{0}}$ is $(\delta_{s,s^{\prime}}-\gamma\cdot\bar{P}_{N_{0}}(s|s^{\prime},a^{\prime}))$. $\bar{\bm{r}}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ with the $(s,a)$-th element being $\bar{r}_{N_{0}}(s,a)$. We have the following result, where the proof is relegated to Section 10.
Lemma 4.1
As long as $\lambda_{N_{0}}=\mathrm{Rad}(N_{0},\varepsilon)$, the following inequality
$V\leq\bar{V}_{N_{0}}+\mathrm{Rad}(N_{0},\varepsilon)\leq V+\mathrm{Rad}(N_{0},\varepsilon)+\frac{2\mathrm{Rad}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(1+\frac{|\mathcal{S}|}{1-\gamma}\right)+\frac{\mathrm{Rad}^{2}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(|\mathcal{S}|+\frac{|\mathcal{S}|^{2}}{1-\gamma}\right).$ (14)
holds with probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\varepsilon$.
Therefore, we can simply set
$\mathrm{Gap}_{1}(N_{0},\varepsilon)=\mathrm{Rad}(N_{0},\varepsilon)$ (15)
and
$\mathrm{Gap}_{2}(N_{0},\varepsilon)=\frac{2\mathrm{Rad}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(1+\frac{|\mathcal{S}|}{1-\gamma}\right)+\frac{\mathrm{Rad}^{2}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(|\mathcal{S}|+\frac{|\mathcal{S}|^{2}}{1-\gamma}\right).$ (16)
5 Algorithm Design
In this section, we present our algorithms for solving the CMDP problem. We first show how to identify one optimal basis of LP (8) and then present the formal algorithm to achieve a logarithmic level regret bound.
5.1 Characterize One Optimal Basis
We now describe how to identify one optimal basis of the LP (8), by sequentially discarding the sub-optimal actions and the redundant constraints. We will use the estimates constructed in the previous Section 4 to carry out the discarding procedure. The formal algorithm to identify such non-zero elements and the constraints is given in Algorithm 1.
Algorithm 1 Algorithm for identifying one optimal basis
1:Input: the historical sample set $\mathcal{F}_{N_{0}}$ that contains $N_{0}$ samples for each $(s,a)\in\mathcal{S}\times\mathcal{A}$.
2:Compute the value of $\bar{V}_{N_{0}}$ as in (10).
3:Initialize $\mathcal{I}$ to be the whole index set that contains every column index of matrix $B$ in (8) and $\mathcal{J}=[K]$.
4:for $i\in\mathcal{I}$ do
5:     Let =\$\𝑖\mathcal{I}^{\prime}=\mathcal{I}\backslash\{i\}$.
6:     Compute the value of $\bar{V}_{\mathcal{I}^{\prime},N_{0}}$ as in (17).
7:     If $|\bar{V}_{\mathcal{I}^{\prime},N_{0}}-\bar{V}_{N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\varepsilon)+2\mathrm{Gap}_{2}(N_{0},\varepsilon)$, then we set $\mathcal{I}=\mathcal{I}^{\prime}$.
8:end for
9:for $k=1,\dots,K$ do
10:     Let $\mathcal{J}^{\prime}=\mathcal{J}\backslash\{q\}$.
11:     Compute the value of $\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}$ as in (20).
12:     If $|\bar{V}_{N_{0}}-\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\varepsilon)+2\mathrm{Gap}_{2}(N_{0},\varepsilon)$, then we set $\mathcal{J}=\mathcal{J}^{\prime}$.
13:end for
14:Output: the set of indexes $\mathcal{I}$ and $\mathcal{J}$.
We now explain the intuition why Algorithm 1 works. Denote by $\bm{q}^{*}$ an optimal solution to LP (8). We describe how to identify the non-zero elements in $\bm{q}^{*}$ and how to identify the constraints such that the values of the non-zero elements of $\bm{q}^{*}$ can be uniquely determined by the corresponding linear equation. For each $i$-th element of $\bm{q}^{*}$, we compare the value of $V$ (8) against $V$ with an additional constraint that $q_{i}=0$. If the two values are different, we identify a non-zero element. To this end, for an index set $\mathcal{I}$, we define an LP, as well as its estimate, as follows.
$V_{\mathcal{I}}=$ $\max$ $\hat{\bm{r}}^{\top}\bm{q}$ $\bar{V}_{\mathcal{I},N_{0}}=$ $\max$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ (17)
    s.t. $C\bm{q}\leq\bm{\alpha}$     s.t. $\bar{C}_{N_{0}}\bm{q}\leq\bm{\alpha}+\lambda_{N_{0}}$
$B\bm{q}=\bm{\mu}$ $|\bar{B}_{N_{0}}\bm{q}-\bm{\mu}|\leq\lambda_{N_{0}}$
$\bm{q}_{\mathcal{I}^{c}}=0$ $\bm{q}_{\mathcal{I}^{c}}=0$
$\bm{q}\geq 0,$ $\bm{q}\geq 0.$
where $\mathcal{I}^{c}$ denotes the complementary set of $\mathcal{I}$. Note that if $V-V_{\mathcal{I}}>0$, we know that $\mathcal{I}^{c}$ contains a non-zero basic variable. The steps 4-8 in Algorithm 1 reflect this point. Starting from $\mathcal{I}$ denoting the whole index set, we sequentially delete one element $i$ (denoting $(s,a)$ in the infinite horizon discounted problem) from the set $\mathcal{I}$. Once we detected that $V-V_{\mathcal{I}\backslash\{i\}}>0$, we know that $i$ is a non-zero basic variable and we add $i$ back into the set $\mathcal{I}$. In this way, we can classify all the basic variables into the set $\mathcal{I}$. However, since we do not know the exact value of $V$ and $V_{\mathcal{I}}$, we use the estimates and compare the value of $\bar{V}_{N_{0}}$ and $\bar{V}_{\mathcal{I},N_{0}}$. For this comparison to be valid, the estimation error has to be smaller than the intrinsic gap between $V$ and $V_{\mathcal{I}}$. We define a constant
$\delta_{1}=\min_{\mathcal{I}}\{V-V_{\mathcal{I}}:V-V_{\mathcal{I}}>0\}$ (18)
and we need $N_{0}$ to be large enough such that the estimation gap is smaller than $\delta_{1}/2$. The quantity $\delta_{1}$ plays a similar role as the gap between the optimal arm and the sub-optimal arm in the multi-arm-bandit problem when deriving a problem-dependent bound (e.g. Lai and Robbins (1985)).
Denote by $M$ the number of non-zero basic variables detected in Algorithm 1. We then denote by $M^{\prime}$ the number of binding constraints corresponding to the optimal solution given by the optimal basis we detected. Our goal is to identify the linearly independent binding constraints such that the non-zero basic variables can be uniquely determined as the solution to a linear equation, as described in Lemma 3.1.
To tell which constraints are redundant and thus can be deleted, we consider the dual program of $V_{\mathcal{I}}$, where $\mathcal{I}$ is determined in steps 4-8 in Algorithm 1, and similarly, we test which dual variable can be set to 00 without influencing the dual objective value. For a dual variable index subset $\mathcal{J}\subset[K]$, we consider the dual program as follows.
$\text{Dual}_{\mathcal{J},\mathcal{I}}=\min\leavevmode\nobreak\ \leavevmode\nobreak$ $\bm{\alpha}^{\top}\bm{y}+\bm{\mu}^{\top}\bm{z}$ (19a)
s.t. $(C(:,\mathcal{I}))^{\top}\bm{y}+(B(:,\mathcal{I}))^{\top}\bm{z}\geq\hat{\bm{r}}_{\mathcal{I}}$ (19b)
$\bm{y}_{\mathcal{J}^{c}}=0$ (19c)
$\bm{y}\geq 0,\bm{z}\geq-\infty,$ (19d)
where $\mathcal{J}^{c}=[K]\backslash\mathcal{J}$. The estimate of $\text{Dual}_{\mathcal{J},\mathcal{I}}$, can be obtained from the estimate of its dual, which is given below.
$\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}=\bar{V}_{\mathcal{J},\mathcal{I},N_{0}}=\max\leavevmode\nobreak\ \leavevmode\nobreak$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ (20a)
s.t. $\bar{C}_{N_{0}}(\mathcal{J},:)\bm{q}\leq\bm{\alpha}+\lambda_{N_{0}}$ (20b)
$|\bar{B}_{N_{0}}\bm{q}-\bm{\mu}|\leq\lambda_{N_{0}}$ (20c)
$\bm{q}_{\mathcal{I}}=0$ (20d)
$\bm{q}\geq 0,$ (20e)
Similarly, we compare the value of $\text{Dual}_{\mathcal{I}}$, where $\text{Dual}_{\mathcal{I}}=\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I}}$ with $\mathcal{J}^{\prime}=[K]$, and $\text{Dual}_{\mathcal{J},\mathcal{I}}$. However, we can only compare the value of their estimates $\text{Dual}_{\mathcal{I},N_{0}}$ and $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$. To this end, we define a constant
$\delta_{2}=\min_{\mathcal{J}\subset[K]}\{\text{Dual}_{\mathcal{J},\mathcal{I}}-\text{Dual}_{\mathcal{I}}:\text{Dual}_{\mathcal{J},\mathcal{I}}-\text{Dual}_{\mathcal{I}}>0\}.$ (21)
For the comparison to be valid, we need $N_{0}$ to be large enough such that the estimation error is smaller than $\delta_{2}/2$. In this way, we identify the linearly independent binding constraints corresponding to $\bm{q}^{*}$, as described in steps 9-13 of Algorithm 1.
5.2 Our Formal Algorithm
We now describe our formal algorithm. From the output of Algorithm 1, we characterize one optimal solution. If the sample size $n$ is used in Algorithm 1, we denote by $\mathcal{I}_{n}$ and $\mathcal{J}_{n}$ the output of Algorithm 1. To be specific, we have $\bm{q}^{*}_{\mathcal{I}_{n}^{c}}=0$ and the non-zero elements $\bm{q}^{*}_{\mathcal{I}_{n}}$ can be given as the solution to
[…]
However, in practice, both the matrices $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ are unknown. We aim to use the samples to learn the matrices $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ such that the $\bm{q}^{*}_{\mathcal{I}_{n}}$ can also be determined. Our formal algorithm is given in Algorithm 2. The steps 3-8 in Algorithm 2 is to use Algorithm 1 as a subroutine to identify the set $\mathcal{I}$ and $\mathcal{J}$ that satisfy the conditions in Lemma 3.1. We can show that as long as $n\geq N_{0}^{\prime}$, where $N_{0}^{\prime}$ is a threshold that depends on the problem parameters, Algorithm 1 correctly obtains the set $\mathcal{I}$ and $\mathcal{J}$ satisfying the conditions in Lemma 3.1. Therefore, we exponentially increase the value of $n$ as input to Algorithm 1 to reach $N_{0}^{\prime}$.
Algorithm 2 The Adaptive-resolving Algorithm
1:Input: the number of samples $N$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$.
2:Initialize $\mathcal{F}_{1}=\emptyset$, $\bm{\alpha}^{1}=N\cdot\bm{\alpha}$ and $\bm{\mu}^{1}=N\cdot\bm{\mu}$.
3:for $n=1,\dots,N$ do
4:     if $n=2^{m}$ for an integer $m$ then
5:         Obtain the output $\mathcal{I}_{n}$ and $\mathcal{J}_{n}$ from Algorithm 1 with the input $\mathcal{F}_{n}$.
6:     else
7:         set $\mathcal{I}_{n}=\mathcal{I}_{n-1}$ and $\mathcal{J}_{n}=\mathcal{J}_{n-1}$.
8:     end if
9:     Construct estimates $\bar{C}^{n}(\mathcal{J}_{n},\mathcal{I}_{n})$ and $\bar{B}^{n}(:,\mathcal{I}_{n})$ using the sample set $\mathcal{F}_{n}$.
10:     Construct a solution $\tilde{\bm{q}}^{n}$ such that $\tilde{\bm{q}}^{n}_{\mathcal{I}_{n}^{c}}=0$ and $\tilde{\bm{q}}^{*}_{\mathcal{I}_{n}}$ is the solution to
[…]
11:     Project $\tilde{\bm{q}}^{n}$ to the set $\{\bm{q}:\|\bm{q}\|_{1}\leq 2\}$ to obtain $\bm{q}^{n}$.
12:     For each $(s,a)\in\mathcal{I}_{n}\subset\mathcal{S}\times\mathcal{A}$, we query the model $\mathcal{M}$ to obtain a sample of the reward $r^{n}(s,a)$ and the costs $c^{n}_{k}(s,a)$ for each $k\in\mathcal{J}_{n}\subset[K]$, as well as the state transition $s^{n}(s,a)$.
13:     Update $\mathcal{F}_{n+1}=\mathcal{F}_{n}\cup\left\{r^{n}(s,a),c^{n}_{k}(s,a),s^{n}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A},\forall k\in[K]\right\}$.
14:     Denote by $\bm{c}^{n}(s,a)=(c^{n}_{k}(s,a))_{\forall k\in\mathcal{J}_{n}}$ and do the update:
$\bm{\alpha}^{n+1}_{\mathcal{J}_{n}}=\bm{\alpha}^{n}_{\mathcal{J}_{n}}-\sum_{(s,a)\in\mathcal{I}_{n}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a).$ (24)
15:     Do the update:
$\bm{\mu}^{n+1}(s^{\prime})=\bm{\mu}^{n}(s^{\prime})-\sum_{(s,a)\in\mathcal{I}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall s^{\prime}\in\mathcal{S}.$ (25)
where $\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}}$ is an indicator function of whether the state transition $s^{n}(s,a)$ equals $s^{\prime}$.
16:end for
17:Output: $\bm{q}^{1},\dots,\bm{q}^{N}$.
A crucial element in Algorithm 2 (step 10) is that we adaptively update the value of $\bm{\alpha}^{n}_{\mathcal{J}_{n-1}}$ and $\bm{\mu}^{n}(s^{\prime})$ as in (24) and (25). We then use the updated $\bm{\alpha}^{n}_{\mathcal{J}_{n-1}}$ and $\bm{\mu}^{n}(s^{\prime})$ to obtain the value of $\tilde{\bm{q}}^{n}_{\mathcal{I}_{n}}$ as in (23). Such an algorithmic design has the following natural interpretation and is the key for us to achieve a logarithmic-level regret. Note that if we adopt a static way to compute $\bm{q}^{n}$, for example, we simply solve (22) with the matrices $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ replaced by their estimates, then the gap between $\bm{q}^{n}$ and $\bm{q}^{*}$ will be at the same order of the estimation error of $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$, which is $O(1/\sqrt{n})$. Thus, the gap between $\frac{\sum_{n=1}^{N}\bm{q}^{n}}{N}$ and $\bm{q}^{*}$ will be at the order of $O(1/\sqrt{N})$, which translates into $O(\sqrt{N})$ regret bound. Instead, we use $\bm{\alpha}^{n}_{\mathcal{J}_{n-1}}$ and $\bm{\mu}^{n}(s^{\prime})$ that are adaptive to the historical information. To be specific, if for one $n\in[N]$ and one binding constraint $k\in\mathcal{J}$, the constraint value under the action $\sum_{n^{\prime}=1}^{n}\bm{q}^{n^{\prime}}/n$ is below the target $\alpha_{k}$, we know that $\alpha^{n}_{k}/(N-n+1)$ is greater than $\alpha_{k}$, which results in a greater constraint value for $\tilde{\bm{q}}^{n+1}$ (it is in fact $\alpha^{n}_{k}/(N-n+1)$ beyond the target $\alpha_{k}$). In this way, any gap between the real-time constraint value and the target will be self-corrected in the next period as we adaptively obtain $\tilde{\bm{q}}^{n+1}$. Therefore, such an adaptive design can re-adjust the possible constraint violation by itself, which results in a regret bound of a lower order. In step 11, we further project $\tilde{\bm{q}}^{n+1}$ to the set $\{\bm{q}\geq 0:\|\bm{q}\|_{1}\leq 2\}$ to obtain $\bm{q}^{n+1}$. This prevents $\tilde{\bm{q}}^{n}$ from behaving ill when $n$ is not large and the estimates of $C(\mathcal{J}_{n},\mathcal{I}_{n})$ and $B(:,\mathcal{I}_{n})$ are not accurate enough. We can show that when $n$ is large enough, $\tilde{\bm{q}}^{n+1}$ automatically belongs to the set $\{\bm{q}\geq 0:\|\bm{q}\|_{1}\leq 2\}$.
Remark. Note that the adaptive resolving step (23) is akin to the idea of resolving an LP in online LP, Bandits with Knapsacks, online allocation problems, where logarithmic or better regret is obtained in a series of work Flajolet and Jaillet (2015), Arlotto and Gurvich (2019), Li and Ye (2022), Li et al. (2021c), Vera and Banerjee (2021), Vera et al. (2021), Bumpensanti and Wang (2020), Jiang et al. (2022). However, since we have identified the optimal basis $\mathcal{I}$ and $\mathcal{J}$ and we stick to this optimal basis to avoid the non-degeneracy assumption, resolving the LP would become resolving the linear equations in (23).
6 Theoretical Analysis
In this section, we conduct our theoretical analysis. The analysis can be divided into two parts. In the first part, we show that Algorithm 1 can successfully help us identify one optimal basis to work with. In the second part, we show how to learn the optimal distribution over the optimal basis we have identified.
6.1 High Probability Bound for Identifying One Optimal Basis
We now present the theorem showing that Algorithm 1 indeed helps us identify one optimal basis with a high probability. In practice, the value of $\varepsilon$ will be set to $1/N$ in the following theorem.
Theorem 6.1
For any $\varepsilon>0$, as long as $N_{0}$ satisfies the condition
$2\mathrm{Gap}_{1}(N_{0},\varepsilon)+2\mathrm{Gap}_{2}(N_{0},\varepsilon)\leq\min\{\delta_{1},\delta_{2}\}$ (26)
the outputs $\mathcal{I}_{N_{0}}$ and $\mathcal{J}_{N_{0}}$ of Algorithm 1 satisfy the conditions described in Lemma 3.1 with probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\varepsilon$. Moreover, the sets $\mathcal{I}_{n}$ and $\mathcal{J}_{n}$ will be common for any $n\geq N_{0}$ satisfying (26), which we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$.
The formal proof of Theorem 6.1 is relegated to Section 11. Note that there can be multiple optimal bases to the LP (8), and our approach essentially identifies one particular optimal basis. To see this point, we sort all the elements of the solution $\bm{q}$, and in Algorithm 1, we try to delete each element $i$ from the set $\mathcal{I}$ from the first one to the last one in this order. For a collection of non-zero basic variables denoted as $(i_{1},\dots,i_{M})$, our approaches identifies the collection with $i_{1}$ being the largest among all collections. Then, among all collections with $i_{1}$ being the first element, our approach identifies the collection with $i_{2}$ being the largest, so on and so forth. Our Algorithm 1 identifies such a collection of non-zero basic variables. Similarly, for the redundant constraints, our approach deletes these constraints from the smallest index in the set $[K]$. Therefore, as long as $n$ (which is $2^{m}$ for an integer $m$) is large enough, our Algorithm 1 outputs the same index sets $\mathcal{I}$ and $\mathcal{J}$ that satisfies the conditions in Lemma 3.1, which we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ from now on.
An important problem parameter related to $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ can be described as follows. Define
$A^{*}=\left[\begin{aligned} &C(\mathcal{J}^{*},\mathcal{I}^{*})\\ &B(:,\mathcal{I}^{*})\end{aligned}\right].$ (27)
We then denote by $\{\sigma_{1}(A^{*}),\dots,\sigma_{|\mathcal{S}|+K^{\prime}}(A^{*})\}$ the eigenvalues of a matrix $A^{*}$. We define $\sigma$ as
$\sigma=\min\left\{|\sigma_{1}(A^{*})|,\dots,|\sigma_{|\mathcal{S}|+K^{\prime}}(A^{*})|\right\}.$ (28)
From the non-singularity of the matrix $A^{*}$, we know that $\sigma>0$.
Remark. It is clear to see that $\sigma$ in (28) relates to the conditional number of the matrix $A^{*}$. In fact, it holds that $\kappa(A^{*})\leq\frac{|\mathcal{S}|+K}{\sigma}$, where $\kappa(A^{*})$ denotes the conditional number of the matrix $A^{*}$. The value of $\sigma$ will show up in our final bounds as it characterizes how hard it is to learn the matrix $A^{*}$. Also, $\sigma$ gives a natural upper bound of the range of the optimal dual variable corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Note that in previous literature that establishes logarithmic regret for online LP/resource allocation (e.g. Vera and Banerjee (2021), Li et al. (2021c), Jiang et al. (2022)), the conditional number of the constraint matrix will also show up in the final bounds. Therefore, we regard the existence of $\sigma$ as the consequence of adopting the online LP/resource allocation framework. On the other hand, for the MDP problem without resource constraints, the matrix $A^{*}$ can simply be represented as the probability transition matrix and one can easily show that $\sigma\geq 1-\gamma$. We will use this bound when deriving our guarantees for MDP problems without resource constraints.
6.2 Regret Bounds
We now show that Algorithm 2 achieves a logarithmic regret, as defined in (3). The next lemma shows that to analyze the regret, it suffices to consider the remaining constraint capacities belonging to the binding constraints. The formal proof is relegated to Section 12.
Lemma 6.2
For the optimal basis identified in Algorithm 1 and the corresponding optimal solution $\bm{q}^{*}$, we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ the output sets as long as $N_{0}$ satisfies the condition (26). We also denote by $(\bm{y}^{*},\bm{c}^{*})$ the corresponding optimal dual solution. Then, it holds that
$N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq\sum_{j\in\mathcal{J}^{*}}y^{*}_{j}\cdot\mathbb{E}[\alpha^{N}_{j}]+\sum_{s\in\mathcal{S}}z^{*}_{s}\cdot\mathbb{E}[\mu^{N}_{s}].$ (29)
Therefore, it suffices to analyze how the “remaining resources” $(\bm{\alpha}_{\mathcal{J}^{*}}^{n},\bm{\mu}^{n})$ behave. We now define
$\tilde{\alpha}_{k}(n)=\frac{\alpha^{n}_{k}}{N-n},\leavevmode\nobreak\ \forall k\in\mathcal{J}^{*}\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }\tilde{\mu}_{s}(n)=\frac{\mu^{n}_{s}}{N-n},\leavevmode\nobreak\ \forall s\in\mathcal{S},\forall n\in[N].$ (30)
The key is to show that the stochastic process $\tilde{\alpha}_{k}(n)$ and $\tilde{\mu}_{s}$ possess some concentration properties such that they will stay within a small neighborhood of their initial value $\alpha^{1}_{k}$ and $\mu^{1}_{s}$ for a sufficiently long time. We denote by $\tau$ the time that one of $\tilde{\alpha}_{k}(n)$ for each $k\in\mathcal{J}^{*}$ and $\tilde{\mu}_{s}$ for each $s\in\mathcal{S}$ escape this neighborhood. Then, both $\text{Regret}_{r}(\pi,N)$ and $\text{Regret}_{k}(\pi,N)$ for each $k\in[K]$ can be upper bounded by $\mathbb{E}[N-\tau]$. From the update rule (24) and (25), we know that
$\tilde{\alpha}_{k}(n+1)=\tilde{\alpha}_{k}(n)-\frac{\sum_{(s,a)\in\mathcal{I}^{*}}\bm{c}^{n}(s,a)\cdot q^{n}(s,a)-\tilde{\alpha}_{k}(n)}{N-n-1},\leavevmode\nobreak\ \forall k\in\mathcal{J}^{*}$ (31)
and
$\tilde{\mu}_{s}(n+1)=\tilde{\mu}_{s}(n)-\frac{\sum_{(s,a)\in\mathcal{I}^{*}}q^{n}(s,a)\cdot(\delta_{s^{\prime},s}-\gamma\mathbbm{1}_{\{s^{\prime}=s^{n}(s,a)\}})-\tilde{\mu}_{s}(n)}{N-n-1}.$ (32)
Ideally, both $\tilde{\alpha}_{k}(n+1)$ and $\tilde{\mu}_{s}(n+1)$ will have the same expectation as $\tilde{\alpha}_{k}(n)$ and $\tilde{\mu}_{s}(n)$ such that they become a martingale. However, this is not true since we have estimation error over $C(\mathcal{J}^{*},\mathcal{I}^{*})$ and $B(:,\mathcal{I}^{*})$, and we only use their estimates to compute $\bm{q}^{n}$. Nevertheless, we can show that $\tilde{\alpha}_{k}(n)$ for each $k\in\mathcal{J}^{*}$ and $\tilde{\mu}_{s}$ for each $s\in\mathcal{S}$ behave as a sub-martingale. Then, from the concentration property of the sub-martingale, we upper bound $\mathbb{E}[\alpha^{N}_{k}]$ for each $k\in\mathcal{J}^{*}$ and $\mathbb{E}[\mu^{N}_{s}]$ for each $s\in\mathcal{S}$. The term $|\mathbb{E}[\alpha^{N}_{k}]|$ for each $k\in[K]\backslash\mathcal{J}^{*}$ can be upper bounded as well. The results are presented in the following theorem. The formal proof is relegated to Section 13 and we provide a brief sketch of the proof for illustration.
Theorem 6.3
Denote by $\bar{\pi}^{N}$ the output policy of Algorithm 2 and denote by $N$ the number of rounds. Then, it holds that
$N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log(N)}{N}\right)$
where the parameters $\alpha=\min_{k\in[K]}\{\alpha_{k}\}$, $\Delta=\min\{\delta^{2}_{1},\delta^{2}_{2}\}$ with $\delta_{1}$ given in (18) and $\delta_{2}$ given in (21), $\sigma$ given in (28). Also, for any $k\in[K]$, we have
$N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\mathbb{E}\left[\bm{q}^{n}\right]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log(N)}{N}\right).$ (33)
Sketch of proof.
We first note that following Theorem 6.1, as long as $n\geq N_{0}$ (or $2N_{0}$ when $N_{0}$ cannot be expressed as the exponent of $2$), we correctly identify the basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ (with a high probability at least $1-O(1/T)$). Then, the resolving step (23) simply becomes
[…]
Our remaining proof can be classified into the following two steps.
  • 1.
    For a fixed $\nu>0$, we define a set
    $\mathcal{X}=\{\bm{\alpha}^{\prime}\in\mathbb{R}^{|\mathcal{J}^{*}|}:\alpha^{\prime}_{k}\in[\alpha_{k}-\nu,\alpha_{k}+\nu],\forall k\in\mathcal{J}^{*}\},$ (35)
    and
    $\mathcal{Y}=\{\bm{\mu}^{\prime}\in\mathbb{R}^{|\mathcal{S}|}:\mu^{\prime}_{s}\in[\mu_{s}-\nu,\mu_{s}+\nu],\forall s\in\mathcal{S}\}.$ (36)
    We can show that $\tilde{\bm{\alpha}}(n)$ and $\tilde{\bm{\mu}}(n)$ behave well as long as they stay in the region $\mathcal{X}$ and $\mathcal{Y}$. To be specific, it automatically satisfies that $\|\tilde{\bm{q}}^{n}_{\mathcal{I}^{*}}\|_{1}\leq 2$ and thus $\bm{q}^{n}_{\mathcal{I}^{*}}$ can be expressed as the solution to the linear equations (34).
    We can also show that $\tilde{\bm{\alpha}}(n)$ and $\tilde{\bm{\mu}}(n)$ stay in the region $\mathcal{X}$ and $\mathcal{Y}$ for a sufficiently long time. To this end, we define a stopping time
    $\tau=\min_{n\in[N]}\{\tilde{\bm{\alpha}}(n)\notin\mathcal{X}\text{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\tilde{\bm{\mu}}(n)\notin\mathcal{Y}\}.$ (37)
    and we derive an upper bound of $\mathbb{E}[N-\tau]$, which scales as $O(\log(N))$.
  • 2.
    By using our upper bound of $\mathbb{E}[N-\tau]$ and combining with Lemma 6.2, we show that the gap over the reward $N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]$ and the gap over the constraint $N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\bm{q}^{n}$ for $k\in\mathcal{J}^{*c}$ can all be upper bounded at the order of $O(\log(N))$. By further exploiting the structure of the matrix $A^{*}$, we can show that the gap over the constraint $N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\bm{q}^{n}$ for $k\in\mathcal{J}^{*}$ can be bounded at the order of $O(\log(N))$ as well.
[…]
6.3 Sample Complexity Bounds
Denote by $\bm{q}^{1},\dots,\bm{q}^{N}$ the output of Algorithm 2. Then, We define $\bar{\bm{q}}^{N}$ such that $\bar{\bm{q}}^{N}_{\mathcal{I}^{*c}}=0$ and
$\bar{\bm{q}}^{N}_{\mathcal{I}^{*}}=\frac{1}{N}\cdot\sum_{n=1}^{N}\bm{q}^{n}_{\mathcal{I}_{N}}.$
Here, $\mathcal{I}^{*}$ can be replaced by $\mathcal{I}_{N}$, the output of Algorithm 1 with the input sample size $N$. From Theorem 6.1, as long as $N\geq N_{0}$ where $N_{0}$ satisfies the condition (26), we know that $\mathcal{I}_{N}=\mathcal{I}^{*}$ with a high probability. We can construct a policy $\bar{\pi}^{N}$ as follows. For each $(s,a)\in\mathcal{S}\times\mathcal{A}$, it holds
$P(\bar{\pi}^{N}(s)=a)=\left\{\begin{aligned} &\frac{\bar{q}^{N}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}(s,a^{\prime})},&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}(s,a^{\prime})=0.\end{aligned}\right.$ (38)
We aim to bound the gap between $V_{r}(\pi^{*},\mu_{1})$ and $V_{r}(\bar{\pi}^{N},\mu_{1})$, as well as $V_{k}(\bar{\pi}^{N},\mu_{1})$ and $\alpha_{k}$ for each $k\in[K]$. Note that in the definition (38), the value of $\bar{\pi}^{N}$ is essentially random, where the randomness comes from the filtration $\mathcal{F}_{N}$ that decides the value of $\bm{q}^{1},\dots,\bm{q}^{N}$. Therefore, in the definition of $V_{r}(\bar{\pi}^{N},\mu_{1})$ and $V_{k}(\bar{\pi}^{N},\mu_{1})$ for each $k\in[K]$, we take expectation over the filtration $\mathcal{F}_{N}$. With the above definitions, the regret bound established in Theorem 6.3 can be transferred into the sample complexity bound of the policy $\bar{\pi}^{N}$ given below. The formal proof is relegated to Section 14.
Theorem 6.4
With a sample complexity bound of
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 (defined in (38)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
Note that one caveat in transferring the bounds in Theorem 6.3 into the sample complexity bound of the policy $\bar{\pi}^{N}$ is that the vector $\bar{\bm{q}}^{N}$ is not an occupancy measure since there is a gap between $B\bar{\bm{q}}^{N}$ and $\bm{\mu}$, though bounded at the order of $O(\log N/N)$. However, one can utilize the structure (22) to show a bound over the gap between $\bar{\bm{q}}^{N}$ and $\bm{q}^{*}$, pointwise. In this way, the gap between their state value functions can be bounded, which induces an additional $O(\log(N))$ term. Therefore, we obtain our final sample complexity bounds, scales as $\frac{O(\log^{2}(1/\varepsilon))}{\varepsilon}$.
Our Algorithm 2 can be directly applied to solving MDP problems without resource constraints and our bounds in Theorem 6.3 and Theorem 6.4 still hold. Note that in the MDP problems, the parameter $\sigma$ can be lower bounded by $1-\gamma$, which follows from the fact that the matrix $A^{*}$ can simply be represented by the probability transition matrix. Then, we have the following sample complexity bound for our Algorithm 2.
Proposition 6.5
For the MDP problems without resource constraints, i.e., $K=0$, with a sample complexity bound of
$O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\xi\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$ (39)
we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 (defined in (38)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
Note that one significant term in our sample complexity bound in Proposition 6.5 is $\xi\cdot\Delta$. Recall that $\xi$ represents the minimal mass of the optimal occupancy measure and $\delta$ represents the change of the objective value of LP (8) if we restrict one non-zero basic variable in an optimal solution to be 00. In this sense, $\xi\cdot\Delta$ can be understood as the gap between the value of the optimal policy and the value of a sub-optimal policy that is restricted from taking the optimal action given a state, corresponding to the instance hardness via corner point as described in Section 3.1. Such a problem-dependent gap shows up broadly in the literature for online decision-making with (logarithmic) problem-dependent guarantees, for example, in the multi-arm-bandit problem (Lai and Robbins 1985) and the reinforcement learning problem (Auer et al. 2008).
In terms of the dependency of our sample complexity bound on other problem parameters such as $|\mathcal{S}|$, $|\mathcal{A}|$, and $1-\gamma$, we compare to the series of work (Sidford et al. 2018, Wainwright 2019, Wang 2020, Agarwal et al. 2020, He et al. 2021), that subsequently achieves a sample complexity bound of $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$, where the dependency over $|\mathcal{S}|$, $|\mathcal{A}|$, and $1-\gamma$ is optimal (Gheshlaghi Azar et al. 2013, Lattimore and Hutter 2012). Our sample complexity bound in (39) has a worse dependency in terms of $|\mathcal{S}|$ and $1-\gamma$. This is because we construct an empirical LP to estimate the value of LP (8) and resolve the linear equation as in (23), where the size of the LP (which is $|\mathcal{S}|$) and the eigenvalues of the matrix $A^{*}$ (which is bounded by $1-\gamma$) will play a part. However, our bound (39) enjoys a better dependency in terms of $\varepsilon$.
7 Extensions
We describe the generalizations of our approach and results in this section. We first describe the generalization to the finite horizon episodic setting, where we maximize the cumulative reward over a finite horizon without discounting. We then describe the generalization to the online learning setting where we can only get the sample of the state we are currently visiting, instead of the sample of an arbitrary state.
7.1 Finite Horizon Episodic Setting
We now generalize our approach and results to the finite horizon episodic setting, where there are $H$ periods in the horizon. Now there are $H$ transition kernels $\{P_{h}\}_{h=1}^{H}$. We let $P_{h}(s^{\prime}|s,a)$ denote the probability of transiting to state $s^{\prime}$ when the state action pair is $(s,a)$ at period $h$. As before, the initial distribution over the state is denoted by $\mu_{1}$. For each period $h$, there is a stochastic reward function $r_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ and a stochastic cost function $c_{h,k}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ for each $k\in[K]$. For each state-action pair $(s,a)$ and each period $h$, we also denote by $\hat{r}_{h}(s,a)=\mathbb{E}[r_{h}(s,a)]$ and $\hat{c}_{h,k}(s,a)=\mathbb{E}[c_{h,k}(s,a)]$. Then, for any Markovian policy $\pi$, the average reward is defined as
$V_{r}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{h=1}^{H}r_{h}(s_{h},a_{h})\mid\mu_{1}\right],$ (40)
and the constraint is defined as
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{h=1}^{H}c_{h,k}(s_{h},a_{h})\mid\mu_{1}\right]\leq\beta_{k},\leavevmode\nobreak\ \forall k\in[K].$ (41)
Note that the policy $\pi$ is allowed to be non-stationary, which means that the behavior of the policy $\pi$ on two different periods can be different, even given the same current state. However, the reward functions $\{r_{h}\}_{h=1}^{H}$, cost functions $\{c_{h,k}\}_{\forall h\in[H],\forall k\in[K]}$, and the transition kernels $\{P_{h}\}_{h=1}^{H}$ are all unknown to the decision maker and needs to be learned from the samples. We still assume the existence of a generative model. {assumption} For each period $h$, each state and action pair $(s,a)$, we can query the model $\mathcal{M}$ to obtain an observation of $r_{h}(s,a)$, $c_{h,k}(s,a)$ for each $k\in[K]$, and the new state $s^{\prime}\in\mathcal{S}$, where the transition from $s$ to $s^{\prime}$ follows the probability kernel $P_{h}(s^{\prime}|s,a)$ independently. We still consider the sample complexity bound of our policy. Denote by $\pi^{*}$ the optimal Markovian policy. For any $\varepsilon$, we aim to find an $\varepsilon$-accurate policy $\pi$ such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\pi,\mu_{1})-\beta_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K],$ (42)
with as few samples as possible.
7.1.1 LP Reformulation and Estimates
For the finite horizon episodic problem, the occupancy measure is defined as $q_{h,\pi}(s,a)$ for each state-action pair $(s,a)$, each period $h\in[H]$, and the policy $\pi$. Now $q_{h,\pi}(s,a)$ represents the probability of taking action $a$ at period $h$ with state $s$, under policy $\pi$. Following a series of work (e.g. Zimin and Neu (2013), Bhattacharya and Kharoufeh (2017), Efroni et al. (2020)), we obtain the following linear programming to represent the optimal policy and the optimal occupancy measure.
$V^{\mathrm{Fi}}=\leavevmode\nobreak\ \max$ $\sum_{h\in[H]}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{r}_{h}(s,a)\cdot q_{h}(s,a)$ (43a)
$\mathrm{s.t.\ }$ $\sum_{h\in[H]}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{c}_{h,k}(s,a)\cdot q_{h}(s,a)\leq\beta_{k}$ $\forall k\in[K]$ (43b)
$\sum_{a\in\mathcal{A}}q_{h}(s,a)=\sum_{s^{\prime},a^{\prime}}P_{h-1}(s|s^{\prime},a^{\prime})\cdot q_{h-1}(s^{\prime},a^{\prime})$ $\forall s\in\mathcal{S},\forall[H]\setminus\{1\}$ (43c)
$\sum_{a\in\mathcal{A}}q_{1}(s,a)=\mu_{1}(s)$ $\forall s\in\mathcal{S}$ (43d)
$q_{h}(s,a)\geq 0$ $\forall h\in[H],\forall s\in\mathcal{S},a\in\mathcal{A},$ (43e)
Again, we can rewrite LP (43) into the standard formulation LP (8). Now, the matrix $B\in\mathbb{R}^{|\mathcal{S}|H\times|\mathcal{S}||\mathcal{A}|H}$. For $h=1$ and $s\in\mathcal{S}$, the element of $B$ at the $s$-th row and $(s,a,1)$-th column is $1$, for $a\in\mathcal{A}$. For $h\geq 2$ and $s\in\mathcal{S}$, the element of $B$ at $(h-1)\cdot|\mathcal{S}|+s$-th row and $(s,a^{\prime},h)$-th column is $1$, while the element of $B$ at $sh$-th row and $(s^{\prime},a^{\prime},h-1)$-th column is $-P_{h-1}(s|s^{\prime},a^{\prime})$, for $s^{\prime}\in\mathcal{S}$ and $a^{\prime}\in\mathcal{A}$. The element of $B$ at other rows and columns is set to be 00. The matrix $C\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{S}|H}$ and the element at $k$-th row and $(s,a,h)$-th column is $\hat{c}_{h,k}(s,a)$, for $k\in[K]$ and for $s\in\mathcal{S}$, $a\in\mathcal{A}$, $h\in[H]$. The vector $\bm{\mu}\in\mathbb{R}^{|\mathcal{S}|H}$ and the first $s$-th element of $\bm{\mu}$ is $\mu_{1}(s)$, for $s\in\mathcal{S}$, and the other elements of $\bm{\mu}$ are 00. Then, in the same way as described in Section 4, we can construct $\bar{V}_{N_{0}}$, $\bar{V}_{\mathcal{I},N_{0}}$, and $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$ with the newly specified matrix $B$ and $C$.
We now specify the choice of $\lambda_{0}$ and bound the estimation gap for the finite horizon episodic setting. To this end, for an optimal basis $\mathcal{I}^{\prime}$ and $\mathcal{J}^{\prime}$ of LP (43), we denote by
$A^{\prime}=\left[\begin{aligned} &C(\mathcal{J}^{\prime},\mathcal{I}^{\prime})\\ &B(:,\mathcal{I}^{\prime})\end{aligned}\right].$ (44)
We then denote by $\{\sigma_{1}(A^{\prime}),\dots,\sigma_{|\mathcal{S}|+K^{\prime}}(A^{\prime})\}$ the eigenvalues of a matrix $A^{\prime}$. We define $\sigma^{\prime}$ as
$\sigma^{\prime}=\min\left\{|\sigma_{1}(A^{\prime})|,\dots,|\sigma_{|\mathcal{S}|+K^{\prime}}(A^{\prime})|\right\}.$ (45)
We have the following lemma as the bound of the estimation error.
Lemma 7.1
As long as $\lambda_{N_{0}}=\mathrm{Rad}(N_{0},\varepsilon)\cdot H$, the following inequality
$V\leq\bar{V}_{N_{0}}+H\cdot\mathrm{Rad}(N_{0},\varepsilon)\leq$ $V+H\cdot\mathrm{Rad}(N_{0},\varepsilon)+\mathrm{Rad}(N_{0},\varepsilon)\cdot\frac{2(K+|\mathcal{S}|\cdot H)}{\sigma^{\prime}}$ (46)
$+\mathrm{Rad}^{2}(N_{0},\varepsilon)\cdot\frac{K\cdot|\mathcal{S}|+|\mathcal{S}|^{2}\cdot H}{\sigma^{\prime}}.$
holds with probability at least $1-H\cdot(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\varepsilon$.
7.1.2 Algorithm and Theoretical Bound
Note that with the LP reformulation $V^{\mathrm{Fi}}$ in (43), our Algorithm 1 and Algorithm 2 apply to the episodic setting. The only change we need to make is to replace the matrix $B$ and $C$ with the one described in Section 7.1.1. Denote by $\bm{q}^{1},\dots,\bm{q}^{N}$ the output of Algorithm 2. Then, We define $\bar{\bm{q}}^{N}$ such that $\bar{\bm{q}}^{N}_{\mathcal{I}^{*c}}=0$ and
$\bar{\bm{q}}^{N}_{\mathcal{I}^{*}}=\frac{1}{N}\cdot\sum_{n=1}^{N}\bm{q}^{n}_{\mathcal{I}_{N}}.$
Here, $\mathcal{I}^{*}$ can be replaced by $\mathcal{I}_{N}$, the output of Algorithm 1 with the input sample size $N$. Then, we can construct a policy $\bar{\pi}^{N}$ as follows. For each $h\in[H]$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, it holds
$P(\bar{\pi}^{N}_{h}(s)=a)=\left\{\begin{aligned} &\frac{\bar{q}^{N}_{h}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})},&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})=0,\end{aligned}\right.$ (47)
where $P(\bar{\pi}^{N}_{h}(s)=a)$ denotes the probability for the policy $\bar{\pi}$ to take the action $a\in\mathcal{A}$ given the state $s\in\mathcal{S}$ at period $h\in[H]$. Similarly, if $\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}_{h}(s,a^{\prime})=0$, we can take an arbitrary action. Then, we have the following bound regarding the sample complexity of the policy $\bar{\pi}^{N}$, as a corollary of Theorem 6.4.
Corollary 7.2
With a sample complexity bound of
$O\left(\frac{(|\mathcal{S}|\cdot H+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\xi\cdot\sigma^{\prime 3}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 (defined in (47)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
7.2 Beyond Generative Model
Note that in the previous sections, we assume the existence of a generative model where we can obtain the sample of reward, cost, and state transition for an arbitrary state and action pair. In this section, we show that our approaches and results can be directly generalized to more involved sampling environments, i.e., the online learning and offline learning environments. We illustrate under the infinite horizon discounted setting described in Section 2. However, our results can also be directly applied to the finite horizon episodic setting described in Section 7.1.
7.2.1 Online Learning
Under the online learning setting, we interact with the true environment to obtain samples. To be specific, suppose that at round $n$ the environment is at state $s_{n}$. Then, in order to obtain the samples, we need to choose an action, denoted by $a_{n}$, and we only obtain the sample of $r(s_{n},a_{n})$, $c_{k}(s_{n},a_{n})$ for each $k\in[K]$, and the state transition $(s_{n},a_{n})\rightarrow s_{n+1}$. This will push the environment to state $s_{n+1}$ and we can only obtain information about state $s_{n+1}$ at the next round. Compared to the generative model described in Section 2 where we can obtain samples for an arbitrary state and action pair, the online learning setting is more restricted in that we can only obtain information about the state we are currently at. The online learning setting is formally described in the following assumption. {assumption} Denote by $\mathcal{M}$ the true environment, with the state space $\mathcal{S}$, action space $\mathcal{A}$, and the transition kernel $P$. Given the current state $s$ and the current action $a$, we obtain a sample of $r(s,a)$, $c_{k}(s,a)$ for each $k\in[K]$, and the state transition $(s,a)\rightarrow s^{\prime}$ following the transition kernel. Then, the state of $\mathcal{M}$ will transit to $s^{\prime}$.
The generalization to the online learning setting is direct. We adopt the idea of off-policy learning, i.e., we use a behavior policy that is different from our policy to generate samples and construct estimates of the underlying CMDP. We then use the obtained samples to form the sample set $\mathcal{F}$ to carry out Algorithm 1 and Algorithm 2. Similar to Li et al. (2021a), we use a stationary policy $\pi_{s}$ as the behavior policy. Given the stationarity of $\pi_{s}$, the sample trajectory can be viewed as a sample path of a time-homogeneous Markov chain over the set of state-action pairs $\mathcal{S}\times\mathcal{A}$. Following Paulin (2015), we make the following uniform ergodicity assumption. {assumption} The Markov chain induced by the stationary behavior policy $\pi_{s}$ is uniformly ergodic. Note that a key property to consider is how often we can obtain a sample of a state-action pair $(s,a)$, which measures the learning efficiency under the stationary behavior policy. Denote by $\mu_{s}$ the stationary distribution over all state-action pairs under the behavioral policy. We then define
$\mu_{\min}=\min_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{s}(s,a).$ (48)
Suppose that we obtain $N$ samples under the stationary distribution $\mu_{s}$. Then, following the standard Hoeffding’s inequality, we know that we obtain the number of samples at least $\frac{\mu_{\min}\cdot N}{2}$ for each state-action pair $(s,a)$, with a high probability $1-O(1/N)$. Therefore, we can just pick the number of samples $\frac{\mu_{\min}\cdot N}{2}$ for each state-action pair $(s,a)$ to construct the data set to serve as input to Algorithm 1 and Algorithm 2. Our algorithms will be the same as the ones under the generative model where we query the model for $\frac{\mu_{\min}\cdot N}{2}$ times. We have the following sample complexity bound under the online learning setting as a corollary of Theorem 6.4.
Corollary 7.3
For any $\varepsilon>0$, with a sample complexity bound of
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\mu_{\min}\cdot\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\mu_{\min}$ is given in (48), $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
7.2.2 Offline Learning
In the offline learning setting, the agent does not interact with the true environment. Instead, the learning task of the agent is conducted using an offline dataset, where it is assumed that the state-action pair follows a reference distribution $\mu_{r}$ (e.g. Rashidinejad et al. (2021), Chen et al. (2022)). Note that in order to implement Algorithm 1 and Algorithm 2, we obtain samples following a uniform distribution $\mu_{u}$ over $\mathcal{S}\times\mathcal{A}$, by querying the generative model. To characterize the distribution shift, we define the following notion of deviation
$D=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left\{\frac{1}{|\mathcal{S}|\cdot|\mathcal{A}|\cdot\mu_{r}(s,a)}\right\}.$ (49)
Then, following the same argument as in the online learning setting, we obtain the following sample complexity bound under the offline learning setting as a corollary of Theorem 6.4.
Corollary 7.4
For any $\varepsilon>0$, with a sample complexity bound of
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot D}{\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $D$ is given in (49), $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
7.3 Numerical Studies
We implement our Algorithm 2 to study the numerical performance. We consider a CMDP problem with the state space $|\mathcal{S}|=10$ and the action space $|\mathcal{A}|=10$. We set the discount factor $\gamma=0.7$. We then randomly generate the probability transition kernel $P$. To be specific, for each state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, and the future state $s^{\prime}\in\mathcal{S}$, we uniformly generate a randomly variable $p_{s,a,s^{\prime}}$. Then, the transition probability is defined as $P(s^{\prime}|s,a)=\frac{p_{s,a,s^{\prime}}}{\sum_{s^{\prime\prime}\in\mathcal{S}}p_{s,a,s^{\prime\prime}}}$. For each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected reward $\hat{r}(s,a)$ is uniformly generated from the interval $[1,2]$ (with the reward for the first action set to be 00). The actual reward $r(s,a)=\hat{r}(s,a)+\eta$, where $\eta$ is uniformly distributed among $[-0.5,0.5]$. There are $K=5$ constraints and for each constraint $k\in[K]$ and each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the expected cost $\hat{c}_{k}(s,a)$ to be uniformly generated from $[1,2]$. The actual cost $c_{k}(s,a)=\hat{c}_{k}(s,a)+\eta^{\prime}$, where $\eta^{\prime}$ is uniformly distributed among $[-0.5,0.5]$.
For each total iterations $N$, We apply Algorithm 2 and obtain the output $\bm{q}^{1},\dots,\bm{q}^{N}$. We compare $\bar{\bm{q}}^{N}$ with the optimal occupancy measure. Since our algorithm is a randomized algorithm, we study the performance of our algorithm in expectation. To be specific, given the problem instance and a fixed $N$, we implement our algorithm repeatedly for $M=500$ rounds. Denote by $\bar{\bm{q}}^{N}_{m}$ the output of our Algorithm 2 at round $m$, for $m\in[M]$. We define the error term as $\text{Err}(N)=\frac{1}{M}\cdot\sum_{m=1}^{M}\|\bar{\bm{q}}_{m}^{N}-\bm{q}^{*}\|_{1}/\|\bm{q}^{*}\|_{1}$. We study how the error term $\text{Err}(N)$ scales with $N$. The results are displayed in Figure 2. As we can see, the error term $\text{Err}(N)$ decreases quickly with respect to $N$. Moreover, since our Algorithm 2 requires only solving a set of linear equations in each iteration, the computation cost of our Algorithm 2 is also moderate.
Figure 2: The computational performance of Algorithm 2. The x-label denotes the size of $N$, while the y-label denotes the error term $\text{Err}(N)$.
8 Concluding Remarks
In this work, we present a new algorithm that achieves logarithmic regrets or $\tilde{O}(1/\varepsilon)$ sample complexity bound for the CMDP problems, for the first time in the literature. To achieve these improvements, we develop a new algorithmic framework to analyze the CMDP problems, motivated by the online LP literature. We consider the LP reformulation of the CMDP problems and we solve the LP in an online manner. Besides introducing the online LP framework, our method is new even compared to the online LP literature, as we do not require the non-degeneracy assumption to derive the logarithmic regret. In our research, we adopt the basic tabular setting to derive our results. We leave the logarithmic regret and $\tilde{O}(1/\varepsilon)$ sample complexity for more involved setting in CMDP problems for future research.